### v2

# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

Abstract
Spin-orbit coupling plays an important role in both fundamental physics and technological applications. Precise estimation of the spin-orbit coupling is necessary for accurate designing across various physical setups such as solid state devices and quantum hardware. Here, we exploit quantum features in a 1D quantum wire for estimating the Rashba spin-orbit coupling with enhanced sensitivity beyond the capability of classical probes. The Heisenberg limited enhanced precision is achieved across a wide range of parameters and does not require fine tuning. Such advantage is directly related to the gap-closing nature of the probe across the entire relevant range of parameters. This provides clear advantage over conventional criticality-based quantum sensors in which quantum enhanced sensitivity can only be achieved through fine-tuning around the phase transition point. We have demonstrated quantum enhanced sensitivity for both single particle and interacting many-body probes. In addition to extending our results to thermal states and the multi-parameter scenario, we have provided an measurement basis to perform close to the ultimate precision.
1 Introduction
The original notion of Spin-Orbit Coupling (SOC) refers to the relativistic quantum mechanical effect, connecting the spin and motional degrees of freedom of electrons in an electric field. SOC plays a crucial role in solid state systems by affecting the energy spectrum in ways that has led to major research areas such as spintronics [1, 2], topological insulators [3, 4], quantum dot arrays [5, 6, 7, 8] and quantum simulation with cold gases under artificial gauge fields [9, 10, 11]. In particular, the Rashba type of SOC [12], which may occur naturally in a solid state material [13] or can be engineered in quantum systems [14, 15, 16], plays a crucial role in a wide range of technological applications. This includes spin field-effect transistors [17], memory and logic devices [18, 19], spin-orbit torque devices [20, 21], spin filters and pumps [22], dissipation-less spin transport [23] and interaction-based quantum transport [24]. In addition, Rashba SOC plays a significant role in emerging quantum technologies from solid state quantum simulators [25, 26, 27, 28] to fault-tolerant quantum computation through formation of Majorana fermions [29] and information encoding in pseudo-spins [30]. Therefore, precise knowledge of the SOC parameters is necessary for studying both equilibrium and non-equilibrium properties of the system. So far, SOC has been measured using various methods such as, electron transport [31, 32], spectroscopy [33, 34], and spin Hall effect [35]. The fundamental and technological importance of SOC and the recent developments in the field of quantum sensing pose a timely question: can the precision, with which the SOC parameters are measured, be enhanced using quantum features?
Quantum sensing has now established itself as a key component of quantum technologies that can surpass the limits of traditional classical sensors [36]. The Cramér-Rao inequality from estimation theory provides a lower bound on the precision of estimating an unknown parameter in terms of inverse of the Fisher information. Generally, as the probe size $L$ is increased, Fisher information grows as $F{\sim}L^{\beta}$. For classical probes, the best achievable scaling is the standard limit where $\beta{=}1$. Meanwhile, exploiting quantum features can improve the precision beyond the capacity of classical probes, with $\beta{>}1$ [37, 36, 38], which is known as quantum-enhanced sensitivity. The concept of quantum-enhanced sensitivity was first introduced in an interferometric setup utilizing Greenberger–Horne–Zeilinger (GHZ) type entangled states [39, 40, 41]. It significantly improves the precision of phase shift measurements to achieve the Heisenberg limit, $\beta{=}2$. Since then, this phenomenon has been verified in various platforms including ion-traps [42], optical systems [43, 44, 45], superconducting qubits [46] and nitrogen-vacancy centers [47]. However, the approach exclusively relies on GHZ-type entangled states, which are challenging to generate and highly susceptible to decoherence and particle loss. Furthermore, this method is limited to a phase generating operation to encode information into the quantum state of the probe. Any disturbance in this unitary process reduces the precision, therefore limits the interferometric sensing method to scale up effectively [48].
Another approach to quantum sensing utilizes many-body probes by leveraging a variety of quantum features. Although interactions have destructive effects in the GHZ-based sensing schemes, they play the primary role in many-body sensors [49]. In particular, quantum criticality has been identified as a resource for achieving quantum enhanced sensitivity. This includes various types of criticalities, such as first-order [50, 51, 52, 53, 54], second-order  [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], Floquet [67, 68], time crystal [69, 70, 71, 72, 73], Stark [74, 75, 76] and quasi-periodic [77] localization, and topological [78, 79, 80] phase transitions. In all these criticalities, a common theme that emerges as the resource for quantum enhancement of the sensing capability is the closing of the energy gap at the phase transition point [49]. Nonetheless, the phase transition occurs only at a particular value of a Hamiltonian parameter which restricts the quantum advantage to a small region around the critical point. As one moves away from the criticality, the energy gap reopens and the quantum enhancement starts to disappear, typically falling back to the standard limit of sensing. This necessitates fine-tuning, making it primarily useful for local sensing where substantial prior knowledge about the parameter of interest is required. Therefore, it is desirable to have a quantum system that feature the gap closing behavior over a wide range of the parameter to be estimated.
In this paper, we show that 1D quantum wires with SOC can provide a platform for Heisenberg precision sensing over a large area in parameter space. This finding provides a route for many-body quantum probe without exploiting only the vicinity of a critical point. Therefore, it opens up a way for Heisenberg precision sensing without fine-tuning. In this work we first consider the single parameter estimation scenario, where the Heisenberg scaling is shown for the cases of a single-particle probe, many-body interacting probe, and thermal probe. We then address the multi-parameter estimation case and finally present the optimal measurement basis.
2 Overview: parameter estimation
To estimate $d$ number of unknown parameters $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})$, the first step is to encode them in a quantum probe whose density operator $\rho_{\bm{\alpha}}$ depends on $\bm{\alpha}$. To determine these parameters, one needs to perform a measurement, which is represented by a set of Positive Operator-Valued Measure (POVM) operators $\{\Pi_{k}\}$ applied to the probe. The outcomes follow a classical probability distribution, where the probability of the $k$-th outcome is given by $p_{k}(\bm{\alpha})=\mathrm{Tr}[\Pi_{k}\rho_{\bm{\alpha}}]$. These outcomes are fed into an estimator function to infer the parameters. For this measurement setup, the uncertainty in estimating $\bm{\alpha}$ can be quantified by the covariance matrix $[\text{Cov}({\bm{\alpha}})]_{i,j}{=}\braket{\alpha_{i}\alpha_{j}}-\braket{\alpha_{i}}\braket{\alpha_{j}}$ Note that the diagonal elements are the single parameter variances $\sigma^{2}_{\alpha_{i}}$. The lower bound in uncertainty is set by the Craémer-Rao inequality [81, 82, 83, 84]
$\text{Cov}({\bm{\alpha}})\geq\frac{1}{M}(\mathcal{F}^{C})^{-1},$ (1)
where $M$ is the number of performed measurements and $\mathcal{F}^{C}$ is the $d\times d$ classical Fisher information matrix (CFIM) with elements
$\mathcal{F}^{C}_{i,j}=\sum_{k}\frac{\partial_{\alpha_{i}}p_{k}(\bm{\alpha})\,\partial_{\alpha_{j}}p_{k}(\bm{\alpha})}{p_{k}(\bm{\alpha})}.$ (2)
This basis-dependent lower bound can be optimized over all possible measurements to give the ultimate obtainable precision in terms of the quantum Cramér-Rao bound [84]
$\text{Cov}({\bm{\alpha}})\geq\frac{1}{M}(\mathcal{F}^{Q})^{-1}.$ (3)
The quantum Fisher information matrix (QFIM) elements can be expressed in terms of the eigen-decomposition of the density operator $\rho_{\bm{\alpha}}{=}\sum_{n,(\lambda_{n}\neq 0)}\lambda_{n}\ket{\lambda_{n}}\bra{\lambda_{n}}$ | in the following way [84],
$\mathcal{F}^{Q}_{i,j}=\sum_{n,(\lambda_{n}\neq 0)}\frac{(\partial_{\alpha_{i}}\lambda_{n})(\partial_{\alpha_{j}}\lambda_{n})}{\lambda_{n}}$
$+\sum_{n,(\lambda_{n}\neq 0)}4\lambda_{n}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}}{\partial_{\alpha_{j}}\lambda_{n}})$
$-\sum_{\begin{subarray}{c}n,m\\ (\lambda_{n},\lambda_{m}\neq 0)\end{subarray}}8\frac{\lambda_{n}\lambda_{m}}{\lambda_{n}+\lambda_{m}}\text{Re}(\braket{\partial_{\alpha_{i}}\lambda_{n}}{\lambda_{m}}\braket{\lambda_{m}}{\partial_{\alpha_{j}}\lambda_{n}}).$ (4)
For a pure state $\rho_{\bm{\alpha}}=\ket{\psi_{\bm{\alpha}}}\bra{\psi_{\bm{\alpha}}}$ |, QFIM becomes
$\mathcal{F}^{Q}_{i,j}=4\text{Re}(\braket{\partial_{\alpha_{i}}\psi}{\partial_{\alpha_{j}}\psi})-\braket{\partial_{\alpha_{i}}\psi}{\psi}\braket{\psi}{\partial_{\alpha_{j}}\psi}).$ (5)
To extract a scalar form from the matrix inequalities, a positive weight matrix $W$ can be multiplied on both sides, followed by a trace operation. In particular, for the choice of $W$ being an identity matrix one gets $\sum_{i}\sigma^{2}_{\alpha_{i}}\geq\text{Tr}((\mathcal{F}^{Q})^{-1})/M$. Another useful choice is to take a $W$ matrix with all elements being zero except $W_{i,i}{=}1$. Then the Cramér-Rao inequality becomes $\sigma^{2}_{\alpha_{i}}\geq(\mathcal{F}^{Q})^{-1}_{i,i}/M\geq 1/M\mathcal{F}^{Q}_{i,i}$, in which the last term is associated with the uncertainty in $\alpha_{i}$ when all the other parameters are known. This situation is known as the single parameter estimation problem where the error in estimating one unknown parameter $\alpha$ is lower bounded by $\sigma^{2}_{\alpha}\geq 1/MF^{C}(\alpha)\geq 1/MF^{Q}(\alpha)$. Here, $F^{C}(\alpha)$ and $F^{Q}(\alpha)$ are the single parameter classical Fisher information (CFI) and quantum Fisher information (QFI), and are just the diagonal elements in CFIM and QFIM, respectively. One choice of optimal basis, in which CFI saturates the QFI, is given by the projcetors to the eigenstates of the so called symmetric logarithmic operator (SLD) $\mathcal{L}$, which are implicitly defined for a probe state $\rho$ as $\partial_{\alpha}\rho{=}(\rho\mathcal{L}{+}\mathcal{L}\rho)/2$.
3 Model: 1D quantum wire
The model considered in this work is based on the transport of electrons in a one-dimensional ballistic quantum wire in the presence of Rashba SOC [85, 86]. Low dimensional transport properties of electrons are of extreme importance in condensed matter physics, with applications in spintronics and quantum information processing. SOC is quite prevalent in two-dimensional electron gas systems with strong confinement along the $z$-axis. Further confinement to a wire geometry enhances the effect of SOC [17] and has led to extensive studies on the transports of non-interacting electrons in quasi 1D systems. The effective dynamics can be described by a tight-binding lattice model, which is governed by the Hamiltonian in the following form
$H=H_{0}+H_{R}+H_{Z}.$ (6)
Here $H_{0}$ describes the spin-independent nearest-neighbor hopping with strength $t$,
$H_{0}=-t\sum_{j,\sigma}(c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma}+c^{{\dagger}}_{j,\sigma}c_{j+1,\sigma}),$ (7)
where $c^{{\dagger}}_{j,\sigma}(c_{j,\sigma})$ is the fermionic creation (annihilation) operator for site $j$ and spin $\sigma$ (denoting $\uparrow$ and $\downarrow$). The Rashba SOC term $H_{R}$ is
$H_{R}=$ $-\alpha_{z}\sum_{j,\sigma,\sigma^{\prime}}(i\sigma_{y})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma^{\prime}}$
[…]
where $\alpha_{y(z)}$ are the SOC parameters and stem from the confining potentials of the wire. The Zeeman term $H_{Z}$ accounts for the presence of an external magnetic field $B$ along z direction,
$H_{Z}=B\sum_{j,\sigma,\sigma^{\prime}}(\sigma_{z})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma^{\prime}}.$ (9)
4 Single Parameter Sensing
In low-dimensional transport experiments, accurate knowledge of the SOC parameters is desired. Therefore, we first focus on estimating the SOC parameter in the uniform Rashba field scenario along $y$- and $z$-direction, i.e. $\alpha_{y}{=}\alpha_{z}$. The non-uniform SOC case will be addressed in the next section. The goal is to analyze how well one can estimate $\alpha_{z}$ using the ground state of the system as a probe. To stay close to the experimental systems, we consider finite wires with $L$ lattice sites and study the growth of QFI with system size. As the scaling of QFI is closely connected to the gap closing of the system, we also analyze the energy gap $\Delta$ between first excited state and ground state. The ground state is doubly degenerate, so a Zeeman term is essential to break the degeneracy. This term need to be strong enough to capture the actual energy gap and not the Zeeman splitting. For example, one should choose $B{>}0.005t$ for $L{=}100$.
4.1 Single particle probe
Figure 1: Scaling analysis of energy gap and QFI. Probe state is the ground state of the Hamiltonian in Eqs. (6)-(9) for sensing $\alpha_{z}$ with hopping parameter $t{=}1$, and $\alpha_{y}{=}\alpha_{z}$. (a) Energy gap between first excited state and ground state $\Delta$ with $\alpha_{z}$ and $B$ for system size $L{=}100$. The four points in the parameter space are chosen as $(B,\alpha_{z})=(0.05,0.1),(0.05,0.9),(0.9,0.9),(0.9,0.05)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling relation of the energy gap with system size at these four parameter points, labeled by corresponding legends. In all cases numerical fit $\Delta{=}aL^{-\mu}{+}b$ shows almost quadratic scaling of the gap closing. (c) QFI of the probe with respect to $\alpha_{z}$ in the parameter space spanned by $\alpha_{z}$ and $B$ for system size $L{=}100$. The same four points are chosen to study the scaling behavior. (d) Numerical fit confirms the quadratic scaling of QFI in the whole range as $F^{Q}{\sim}L^{\beta}$ with $\beta{\approx}2$.
We start the sensing analysis with the single particle ground states of the Hamiltonian in Eq.(6) as the probe. In Fig. 1, we show the behavior of the energy gap $\Delta$ and QFI $F^{Q}$ of the ground state with respect to $\alpha_{z}$. We have considered experimentally relevant values of $\alpha_{z}$ and $B$ for our analysis. Fig 1(a) shows the result for a fixed size $L=100$ where, for weak magnetic fields, the gap goes through a minima at a small value of $\alpha_{z}$. As the field becomes stronger, the minima becomes less pronounced and the gap increases almost monotonically with $\alpha_{z}$. Interestingly, the scaling of the decline of energy gap with system size shows similar behavior across the entire range of $\alpha_{z}$ and $B$. This is shown in Fig. 1(b), where four different combinations of $\alpha_{z}$ and $B$ values in the weak and strong regimes have been chosen to analyze the scaling. In all the cases, the gap closing was found to scale almost quadratically with system size, i.e. $\Delta\sim L^{-\mu}$ with $\mu\approx 2$. Now for the QFI, Fig. 1(c) shows almost a complementary picture to Fig. 1(a), illustrating the close connection between the gap closing and the QFI. This leads to a quadratic scaling of QFI in the whole region spanned by $\alpha_{z}$ and $B$, i.e. $F^{Q}\sim L^{\beta}$ with $\beta\approx 2$, see Fig. 1(d). This is a very promising result as unlike criticality-based quantum sensors, where quantum-enhanced sensitivity is only achievable at the vicinity of the phase transition, one can reach Heisenberg scaling over a wide range of parameters. This can be explained by the favorable gap closing nature of the system. We also note that for a given system size, the sensing capacity is stronger for smaller values of SOC parameters in the presence of a weak external magnetic field.
4.2 Many-body interacting probe
Now we study the effects of two-body contact interactions on the sensing capability of the probe. In one-dimensional systems, such interactions are known to strongly influence the low-energy physics of many-body systems, leading to Luttinger liquid behavior. The Hamiltonian in Eq. (6) is modified to include the interaction between two fermions with different spins at the same site with strength $U$ and between two fermions on adjacent sites with strength $V$. This results in rewriting the Hamiltonian as
$H=H_{0}+H_{R}+H_{Z}+H_{\rm int},$ (10)
with,
$H_{\rm int}=U\sum_{j}c^{{\dagger}}_{j,\uparrow}c_{j,\uparrow}+V\sum_{j,\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma}c^{{\dagger}}_{j+1,\sigma^{\prime}}c_{j+1,\sigma^{\prime}}.$ (11)
The $U$ term accounts for an increase in the system’s energy when two fermions with opposite spins occupy the same site. For fermions in optical lattices, this term can be tuned by changing the scattering length with Feshbach resonance [87]. The $V$ term can arise due to the Coulomb repulsion, for example, between electrons confined in quantum dot arrays. As a typical scenario in theoretical and experimental situations, we consider the half-filled case here, i.e. $L$ fermions on a $L$-site wire. When $U{=}V{=}0$, the fermionic ground state is given by the antisymmetric Slater determinant state formed by the single particle energy eigenstates {|$\{\ket{E_{n}}\}$, which can be compactly written with the symmetric group $S_{L}$ as
|$\ket{\psi_{\rm GS}}=\frac{1}{\sqrt{L!}}\sum_{\bm{n}\in{S_{L}}}\text{sgn}(\bm{n})\ket{E_{n_{1}}}\dots\ket{E_{n_{L}}}.$ (12)
This results in an analytical QFI expression [78],
$F^{Q}=4\left(\sum_{l=1}^{L}\braket{\partial E_{l}}{\partial E_{l}}-\sum_{l,l^{\prime}}\braket{\partial E_{l}}{E_{l^{\prime}}}\braket{E_{l^{\prime}}}{\partial E_{l}}\right).$ (13)
Figure 2: Many-body probe. Probe state is the half-filled ground state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, and $\alpha_{y}{=}\alpha_{z}$. (a) Scaling of energy gap in the non-interacting case ($U{=}V{=}0$) with system size $L$ for different values of SOC parameter $\alpha_{z}$. (b) Scaling of QFI with respect to $\alpha_{z}$ in the non-interacting case for different $\alpha_{z}$ values. (c) QFI at $\alpha_{z}{=}0.1t$ as a function of $U$ when $V{=}0$. (d) QFI scaling at $\alpha_{z}{=}0.1t$ for different $U$ values. (e) Spectrum $E$ as a function of $U$. (f) QFI at $\alpha_{z}{=}0.1t$ as a function of $V$ when $U{=}0$.
Top row of Fig. 2 shows the results of the many-body non-interacting probe where the external field is fixed at $B{=}0.01t$. As the single particle eigenstates are filled up to $L$-th level in the ground state, the energy gap is defined by the energy difference between the $L{+}1$ and $L$-th eigenvalues. As Fig. 2(a) shows, gap closing is still algebraic with system size $(\Delta{\sim}L^{-\mu})$, but it is less than quadratic. The exponent $\mu$ is found to be decreasing with increasing $\alpha_{z}$. On the other hand, the QFI still scales quadratically with system size, as shown in Fig. 2(b). This shows that, the fermionic statistics in the non-interacting many-body ground state still results in the Heisenberg limit of precision for a wide range of the SOC parameter. When only $U$ is non-zero, we observe that the QFI starts increasing with $U$ before attaining a saturation in the repulsive case. For attractive $U$, the QFI decreases and goes to zero. Fig 2(c) displays this behavior, while Fig. 2(d) shows that algebraic scaling of QFI is still sustained in the repulsive case. The exponential growth of the Hilbert space dimension limits our calculations in terms of system size, which can therefore be affected by finite size effects. Nevertheless, the quadratic scaling of QFI for $U{=}0$ is expectedly observed in Fig. 2(d). Surprisingly, as $U$ is increased, the decrease in exponent $\beta$ is not drastic, which indicates that the interacting probes can still provide quantum enhancement. This can be attributed to the fact that the presence of $U$ in Eq. (11) does not open up the ground state energy gap as shown in the band structure in Fig. 2(e). The fermions in the ground state stay localized at different sites. In the other case, when $V$ is non-zero and $U{=}0$, the QFI steadily decreases (see Fig. 2(f)). We numerically observe that quadratic scaling of QFI survives for small values of $V$ and disappears near $V{\approx}t$.
4.3 Thermal probe
Figure 3: Thermal probe. Probe state is the thermal state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, $\alpha_{y}{=}\alpha_{z}$. (a) QFI vs. temperature in the non-interacting case for different system sizes. As thermal energy surpasses the gap, QFI falls of universally as ${\sim}1/T$. (b) Variation of QFI with both $U$ and $T$ for $L{=}6$.
So far we have discussed the cases where the probe is taken to be the ground state which is the equilibrium state at zero temperature. In practice however, the sensors generally operate at finite temperature and hence, it is necessary to analyze the effect of temperature on the sensing capability. When the temperature is smaller than the energy gap i.e. $k_{B}T{<}\Delta$ ($k_{B}$ is the Boltzmann constant), we expect the ground state description to hold. Higher temperatures result in a mixed state with increasing contributions from higher energy states. The QFI then starts to decrease and typically collapses universally with an algebraic decay with temperature [74]. In Fig. 3(a), we show this behavior in the non-interacting case for two system sizes $L=100,200$. The ground state QFI value is retained up to a temperature $T{\sim}\Delta/k_{B}$, after which the QFI is found to fall off as $T^{-1}$. In the repulsive interacting case, we observe that the energy gap decreases with $U$, and hence the temperature at which the QFI starts to fall decreases with $U$ (see Fig. 3(b)).
5 Multi-parameter sensing
Figure 4: Multi-parameter sensing. Probe state is the single particle ground state of the Hamiltonian in Eq. (6) with $t{=}1$, and $B{=}0.01t$. (a) The $(\alpha_{z},\alpha_{z})$ component of the QFIM $\mathcal{F}^{Q}$ across a range of $\alpha_{z}$ and $\alpha_{z}$ for $L{=}100$. Four points chosen in the parameter space for scaling analysis are $(\alpha_{y},\alpha_{z})=(0.05,0.1),(0.9,0.05),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling of $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ at the four points. (c) The $(\alpha_{y},\alpha_{y})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$ at the four points. (e) The $(\alpha_{y},\alpha_{z})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{z}}$ at the four points.
In this section, we consider a more general scenario where we do not presume prior knowledge of the SOC parameters i.e. the Rashba fields in any direction. The goal is to estimate both $\alpha_{y}$ and $\alpha_{z}$ simultaneously, making the problem multi-parametric. To analyze the system’s sensitivity to changes in these parameters, we compute the multi-parameter QFIM $\mathcal{F}^{Q}$, which provides a comprehensive description of the ultimate precision limits for estimating multiple parameters concurrently. In Fig. 4(a) we shows the component $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ as a function of the two parameters. Fig 4(b) shows that $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ reaches super-linear scaling in a wide range of $\alpha_{z}$ and $\alpha_{y}$ and can also attain close to Heisenberg scaling. Similar observation also holds for the component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{y}}$, as shown in Figs. 4(c) and (d). From this analysis we can also infer that the quantum enhancement of the sensing of $\alpha_{y}$ when $\alpha_{z}$ is known, can reach the Heisenberg limit as well. In general, however, the sensitivity depends on the off-diagonal component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ that captures the correlation between the estimates of $\alpha_{z}$ and $\alpha_{y}$. Fig 4(e) and (f) show that the scaling of $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ also stays close to quadratic across the whole range.
6 Optimal measurement
Figure 5: Comparison between QFI and CFI. (a) For a single-particle probe with $L{=}100$, the top curves correspond to ground state and bottom curves correspond to thermal state at $k_{B}T{=}0.01t$. (b) Many-body interacting probe with $L{=}6$ and $U{=}5t$.
While the QFI gives the ultimate bound for precision, the optimal measurement basis is not unique. In our work, we find that the lattice version of the particle current operator [88, 89], defined as $\mathcal{I}=i\sum_{j,\sigma}(c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma}-c^{{\dagger}}_{j,\sigma}c_{j+1,\sigma})$, generates a suitable basis for estimation. The CFI measured in the eigenbasis of $\mathcal{I}$ matches with the corresponding QFI for single-particle probe, as shown in Fig. 5(a). Thermal energy exceeding the gap and interaction result in small discrepancy between the CFI and QFI for small $\alpha_{z}$ , as shown in Figs. 5 (a) and (b), respectively.
7 Conclusion
Due to the long reaching impact of SOC on various fields ranging from condensed matter systems to emerging quantum technologies, it is of utmost importance to estimate the SOC strength with high precision. In this paper, we employ tools from quantum sensing to achieve quantum enhanced sensitivity for the estimation of Rashba SOC in one dimensional quantum wires. This enhancement is closely connected to gap closing feature of the system. Unlike conventional criticality based quantum sensors, in which the energy gap closing happens at the phase transition point, our probe achieves quantum advantage with Heisenberg precision across a wide range. Our analysis with both single-particle and many-body interacting probes establishes quantum advantage for all such parameters within characteristic temperature range. We have extended our results to multi-parameter sensing scheme where multiple SOC terms can be jointly estimated. In addition, we show that a simple particle current measurement can closely reach the ultimate precision limit.
YB acknowledges support from National Natural Science Foundation of China (Grant No. 12404551) and the China Postdoctoral Science Foundation (Grant No. 2024M750339). AB acknowledges support from National Natural Science Foundation of China (Grants Nos. 12050410253, 92065115, and 12274059) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012).